OpenStudy (anonymous):

x – 3 ≤ – 12 = ? hheeeellpp

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

Yes!c:

OpenStudy (anonymous):

Okay. What would we do to the Left Hand Side in order to isolate x (in other words get rid of the minus 3)?

OpenStudy (anonymous):



OpenStudy (anonymous):

Add 3? right?

OpenStudy (anonymous):

yes and @some_someone seems to have done the operation for you

OpenStudy (anonymous):

yeah to both sides.

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

the trick I follow for negative values if to rewrite it for example -12+3=3-12=-9 or if they are both negatives you can factor out a negative one example -12-3=-1*(3+12)=-1*15=15